Find the mean and standard deviation for the data. Thanks!particular city gave the following statistics: Xi-11,000, and xi 2-1,900,000.
Answer #1

Given:

n = 100

\sum x i = 11000

\sum x i2 = 1900000

Mean = \sum x i / n =  11000 /100

= 110

Standard Deviation =  

= \sqrt{}{1900000 - 121000000/100 } /100

= 83.06